find three consecutive odd integers such that three times the middle integer is one more than the sum of the first and the third
Answer by mananth(16949)

let the 3 consequtive odd integers be x,x+2,x+4
3(x+2)= x+x+4 +1
3x+6 = 2x+5
3x-2x=5-6
x=-1
the integers are -1,1,3